A cook knocks a 95.0 g egg from a kitchen counter top. When the egg is 65.0 cm above the ground, it has a total mechanical energy of 1.12 J. Determine the kinetic energy of the egg.
1
Expert's answer
2020-12-23T07:34:47-0500

The total mechanical energy consists of potential and kinetic energies:


"E_M=E_P+E_K=mgh+mv^2/2,\\\space\\
E_K=E_M-E_P=E_M-mgh,\\
E_K=1.12-0.095\cdot9.8\cdot0.65=0.51\text{ J}."